OpenStudy (anonymous):

What is the refractive index of a medium when the light strikes the medium after traveling through air at an angle of 23.3° and the refracted angle is 14.6°?

OpenStudy (aaronq):

\(\dfrac{\eta_1}{\eta_2}=\dfrac{sin\theta_2}{sin\theta_1}; \eta\)= refractive index, \(\theta \) =angle refraction index,\(\eta\), of air is 1. http://hyperphysics.phy-astr.gsu.edu/hbase/geoopt/refr.html http://hyperphysics.phy-astr.gsu.edu/hbase/tables/indrf.html